Question 5: (15 points). (Bond valuation relationships) Arizona Public Utilities issued a bond that pays $70 in interest, with a $1,000 par value and matures in 25 years. The markers required yield to maturity on a comparable-risk bond is 8 percent. (Round to the nearest cent.) For questions with two answer options (e.g. increase/decrease) choose the best answer and write it in the answer block. Question a. What is the value of the bond if the markers required yield to maturity on a comparable-risk bond is 8 percent? Answer b. What is the value of the bond if the markers required yield to maturity on a comparable-risk bond increases to 11 percent? C. What is the value of the bond if the markets required yield to maturity on a comparable-risk bond decreases to 7 percent? d. The change in the value of a bond caused by changing interest rates is called interest-rate risk. Based on the answer. in parts b and c, a decrease in interest rates (the yield to maturity) will cause the value of a bond to (increase/decrease By contrast, an increase in interest rates will cause the value to (increase/decrease) Also, based on the answers in part b, if the yield to maturity (current interest rate) equals the coupon interest rate, the bond will sell at (parface value) exceeds the bonds coupon rate, the bond will sell at a (discount/premium) and is less than the bonds coupon rate, the bond will sell at a (discount/premium) e. Assume the bond matures in 5 years instead of 25 years, what is the value of the bond if the yield to maturity on a comparable-risk bond is 8 percent? $ 960.07 Assume the bond matures in 5 years instead of 25 years, what is the value of the bond if the yield to maturity on a comparable-risk bond is 11 percent? f. Assume the bond matures in 5 years instead of 25 years, what is the value of the bond if the ield to maturity on a comparable-risk bond is 7 percent? From the findings in part e, we can conclude that a bondholder owning a long-term bond is exposed to (more/less) interest-rate risk than one owning a short-term bond.
Answer #1

As per rules I am answering the first 4 subparts of the question

A: Using financial calculator
Input:

FV= 1000,

PMT=70

N=25

I/Y =8

Find PV = - 893.25

Hence price of the bond is $893.25

B: FV= 1000,

PMT=70

N=25

I/Y =11

Find PV = - 663.13

Hence price of the bond is 663.13

C:

FV= 1000,

PMT=70

N=25

I/Y =7

Find PV = - 1000

Hence price of the bond is 1000

D: Decrease in interest rates will cause the value of bond to increase.

Increase in interest rates will cause the value of the bond to decrease.

If the YTM equals coupon rate, the bond will sell at par

Exceeds the bond rate, the bond will sell at discount

Less than coupon rate, the bond will sell at a premium.